HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INTERNAL ROTATION IN 4-FLUOROACETOPHENONE AND METHYL

4-FLUOROBENZOATE AND THEIR MONO-HYDRATES

\underline{X. WANG}, J.-U. GRABOW, Institut für Physikalische Chemie und Elektrochemie, Leibniz Universität Hannover, 30167 Hannover, Germany ; J. LI, Y. ZHENG, Q. GOU, School of Chemistry and Chemical Engineering, Chongqing University,

Daxuecheng South Rd. 55, 401331, Chongqing

The rotational spectra of 4-Fluoroactophenone and Methyl 4-fluorbenzoate have been investigated by rotational spectroscopy, using a Fourier transform microwave spectrometer operated in the frequency ranges 2-20 GHz^{1} and 2-26.5 GHz^{2}, respectively. Quantum chemical calculations were employed to perform conformational analyses, obtaining only one stable conformer at the MP2/6-311+G(d,p) level for both molecules. Arising from the different CH_{3}O- and CH_{3}- groups of 4-Fluoroactophenone and Methyl 4-fluorbenzoate, the changes in the structure and the V_{3} barrier of the methyl internal rotation were analyzed. Furthermore, the non-covalent interactions in 4-Fluoroactophenone-H_{2}O and Methyl 4-fluorbenzoate-H_{2}O were visualized using Bader’s quantum theory of atoms-in-molecules^{3} and Johnson’s non-covalent interaction analysis^{4}. The rotational spectrum of 4-Fluoroactophenone-H_{2}O has already been analyzed. Interestingly, the same non-covalent interaction binding sites are also present in Methyl 4-fluorbenzoate-H_{2}O, but the preference in the binding sites for the global minimum reverses between the two most stable isomers of 4-Fluoroactophenone-H_{2}O and Methyl 4-fluorbenzoate-H_{2}O at the MP2/6-311++G(d,p) level. The molecular electrostatic potential^{5} of 4-Fluoroactophenone and Methyl 4fluorbenzoate was mapped to visualize the different effect of the CH_{3}O- and CH_{3}- groups on the electron densities, which explains why the preference for the global minimum is predicted to be opposite.

1doi:10.1063/1.4994865, J. Chen, Y. Zheng, J. Wang, G, Feng, Z. Xia and Q. Gou, J. Chem. Phys., 147, 094301, (2017).

2doi:10.1016/j.jms.2012.07.006, M. K. Jahn, D. A. Dewald, D. Wachsmuth, J.-U. Grabow and S. C.

Mehrotra, J. Mol. Spectrosc., 280, 54-60, (2012).

3doi:10.1021/cr00005a013, R. F. W. Bader, Chem. Rev., 91, 893-928, (1991).

4doi:10.1021/ja100936w, E. R. Johnson, S. Keinan, P. Mori-Sánchez, J. Contreras-García, A. J.

Cohen and W. Yang, J. Am. Chem. Soc., 132, 6498-6506, (2010).

5doi:10.1039/D1CP02805G, J. Zhang and T. Lu, Phys. Chem. Chem. Phys., 23, 20323-20328, (2021).

p-number: p128  Submitted on Wed, 14 Jun 2023 12:31:23 +0200